[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Thu Jul 13, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Thu Jul 13 20:52:02 UTC 2017


HSU Shift Log for Thu Jul 13 2017 16:51:57 GMT-0400 (EDT)

Shift: H

Forecaster: J Beven

Technical Issues Summary:

ATCF issues: None.

N-AWIPS issues: None.

AWIPS-2/NCP issues: None.

Other Techincal issues: GMI data not displaying properly on NRL/FMNOC  
center pages. It is available through the NRL directory.

Miscellaneous issues: None.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170713/70cb08fe/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list